# An Empirical Study on LLM-based Agents for Automated Bug Fixing

### v1

@@ Abstract.
Large language models (LLMs) and LLM-based Agents have been applied to fix bugs automatically, demonstrating the capability in addressing software defects by engaging in development environment interaction, iterative validation and code modification. However, systematic analysis of these agent and non-agent systems remain limited, particularly regarding performance variations among top-performing ones. In this paper, we examine seven proprietary and open-source systems on the SWE-bench Lite benchmark for automated bug fixing. We first assess each system’s overall performance, noting instances solvable by all or none of these sytems, and explore why some instances are uniquely solved by specific system types. We also compare fault localization accuracy at file and line levels and evaluate bug reproduction capabilities, identifying instances solvable only through dynamic reproduction. Through analysis, we concluded that further optimization is needed in both the LLM itself and the design of Agentic flow to improve the effectiveness of the Agent in bug fixing.
@@ 1. Introduction
Large Language Models (LLMs) (Zhao et al., 2023) are advanced machine learning models trained on vast amounts of textual data, capable of understanding and generating human-like text. LLM-based Agents (Xi et al., 2023) are systems that utilize large language models to interact with the environment and accomplish specific tasks. Recently, LLM-based Agents have demonstrated significant influence in automated bug fixing in code repositories (Kang et al., 2023; Zhang et al., 2024; Yang et al., 2024a). Thanks to the powerful natural language processing capabilities of LLMs, these Agents can efficiently understand and analyze source code and its associated natural language descriptions, such as user-submitted issue descriptions and code comments. Additionally, through dynamic interaction with local environments (e.g., via terminal), LLM-based Agents can retrieve useful information from the code repository, perform code editing and execution, and iterate and validate repair results, thereby improving the accuracy and efficiency of bug fixes. This combination of LLM and environmental feedback has made automated bug fixing more efficient and feasible than ever before, providing revolutionary new tools for software maintenance and development.
@@
Researchers from both the industry (Liu et al., 2024a; hon, [n. d.]; gru, [n. d.]; Ma et al., 2024) and academia (Zhang et al., 2024; Tao et al., 2024) have developed LLM-based Agent systems to locate and fix bugs in code repositories. To evaluate the fault localization and repair capabilities of LLMs and various Agent systems, Jimenez et.al (Jimenez et al., 2023) proposed the evaluation datasets SWE-bench, with derived versions SWE-bench Lite (a subset of the full benchmark), and SWE-bench Verified (human annotated subset of SWE-bench published recently). These datasets contain real bugs from code repositories and can verify the correctness of the patches generated by Agents through unit tests. Recently, these datasets have become the most influential benchmarks in the field of automated bug fixing, attracting both academic and industrial participants to compete on the SWE-bench Lite leaderboard  [^1] https://www.swebench.com/, with new submissions typically every one or a half week.
@@
However, no work has systematically analyzed the fault localization and repair capabilities of LLM-based Agents or the performance differences among various tools within these Agent systems. Regarding the SWE-bench Lite dataset itself, due to the quality of issue descriptions and the complexity of the logical dependencies related to the defects, some instances in the benchmark are easier for Agents to fix, while others are more difficult (Xia et al., 2024). As for the design of the systems, different designs exhibit different planning, reasoning, and problem solving capabilities, i.e. some systems adopting static approaches (gru, [n. d.]) and others adopting dynamic approaches (Zhang et al., 2024). We have also observed significant differences in the sets of cases that each system can solve. Therefore, analyzing the solving capabilities of LLM-based Agents on specific instances can not only help us better understand the current performance of Agents but also provide comparative insights to inspire future research directions.
@@
We collected the four most outstanding commercial systems (i.e. MarsCode Agent (Liu et al., 2024a), Honeycomb (hon, [n. d.]), Gru (gru, [n. d.]), Alibaba Lingma Agent (Ma et al., 2024)) and the three most excellent open-source systems (i.e. AutoCodeRover (Zhang et al., 2024), Agentless + RepoGraph (Ouyang et al., 2024), Agentless (Xia et al., 2024)) with top performances on the SWE-bench Lite Leaderboard and conducted a comprehensive analysis of the performance differences of each system. First, we evaluated the overall performance of LLM-based Agents in bug fixing, including statistics on the instances that all seven systems can solve and those that none can solve, and analyzed the reasons behind these results. We also explored why some instances can only be solved by Agent systems while others can only be solved by non-Agent systems. Next, we investigated the performance differences in fault localization among different systems and their causes, compiling file-level and line-level localization accuracy rates. Finally, we analyzed the impact of bug reproduction on bug fixing, and the common characteristics of instances that can only be solved through dynamic reproduction.
@@
Through data analysis, we have summarized several insights.To improve bug fixing, it is essential to enhance the model’s reasoning ability, enabling it to accurately identify bug-related information within an issue and reduce noise interference. For multiple potential repair locations, the model should leverage its reasoning capabilities to determine the most relevant location. From the Agentic flow perspective, there should be a strong focus on the quality of the issue and attention to multiple suspicious locations in the stack trace. The design should include mechanisms to verify the completeness of patches and consider their global impact. Mechanisms should also be implemented to mitigate the randomness of the model’s output or effectively utilize its diversity. In error localization, achieving line-level accuracy is more critical than file-level, due to the larger discovery space, necessitating finer-grained results. During the reproduction process, ensuring the correctness of reproductions is crucial, as incorrect reproductions can result in the failure of the entire solving process.
Novelty and Contribution To the best of knowledge, this is the first work to:
  1. (1)
@@
    Study the effectiveness of LLM-based Agents in automatic bug fixing for code repositories
  2. (2)
@@
    Examine the effectiveness of different LLM-based Agents in Fault Localization and analyze the reasons for their differences
  3. (3)
    Investigate the impact of bug reproduction on bug fixing of LLM-based Agents
  4. (4)
@@
    Summarize the current issues and future research directions for LLM-based Agents in bug fixing
@@ 2. Background
In this section, we first introduce SWE-bench Lite and then we introduce the leading LLM-based bug fixing systems.
2.1. SWE-bench Lite
@@
As the computational demands and high difficulty of SWE-bench which comprises 2,294 issue-commit pairs across 12 Python repositories, the authors of SWE-bench introduces SWE-bench Lite  [^2] https://www.swebench.com/lite.html, which includes 300 more manageable and self-contained instances focused on functional bug fixes, covering 11 of the original 12 repositories. It retains the diversity of SWE-bench but is easier to evaluate. The selection criteria is shown below:
  1. (1)
    Removed instances with images, external links, commit SHAs, or references.
  2. (2)
    Excluded problem statements under 40 words.
  3. (3)
    Excluded instances editing more than one file or with more than three edit hunks.
  4. (4)
    Excluded instances creating/removing files or with error message checks.
  5. (5)
    Sampled final 300 test and 23 development instances from remaining ones.
@@ 2.2. Leading LLM-based Bug Fixing Systems
LLM-based Bug Fixing Systems are systems built on Large Language Models (LLMs) that can automatically edit code repositories to fix bugs based on issue reports. Bug fixing is a highly resource-intensive task in software development, requiring developers to reproduce the bugs reported in issue reports, precisely locate defective code snippets within large code repositories, understand the cause of errors, and implement fixes. Automating bug fixing has long attracted widespread attention in both academia and industry. Given the strong logical reasoning and coding capabilities demonstrated by LLMs, numerous works have explored the development of automated bug fixing tools based on LLMs. In this paper, we study seven leading LLM-based Bug Fixing Systems (four commercial systems (Liu et al., 2024a; hon, [n. d.]; gru, [n. d.]; Ma et al., 2024) and three open source systems (Zhang et al., 2024; Xia et al., 2024; Ouyang et al., 2024)), comparing their differences in system design and performance in automated bug fixing, analyzing the shortcomings and limitations of existing systems, and providing direction for future work in building adaptive, high-reliability automated bug fixing systems.
Table  1 shows the leading LLM-based bug fixing systems on SWE-bench Lite. We classified the fault localization techniques adopted by different systems from three perspectives: (1) the utilization of LLMs, (2) the utilization of CodeGraph, and (3) the utilization of software analysis techniques. We also compared whether different systems use reproduction for patch validation and analyzed the patch generation format. Blow this section, we will introduce these systems from a technical perspective:
MarsCode Agent (Liu et al., 2024a),developed by ByteDance, is a bug fixing system that combines code knowledge graphs, software analysis techniques, and LLMs. It uses a Reproducer Agent to automatically reproduce the bugs described in the issue reports. The system constructs a code knowledge graph for the code repository and utilizes graph reasoning, software analysis techniques, and the reasoning capabilities of LLMs to achieve fine-grained defect localization. Marscode Agent generates candidate patches using LLMs and selects the final patch from these candidates based on the reproduction scripts. It has demonstrated outstanding performance in bug fixing tasks, achieving the highest issue resolution rate on the SWE-bench Lite leaderboard.
Honeycomb (hon, [n. d.]), an agent-based bug fixing system developed by Honeycomb. It provides tools for file viewing and editing to the LLM, enabling it to attempt to construct reproduction scripts for all issues. The system calls these tools for defect localization and generates patches in the format of $(Line_{b},Line_{e},Replace)$. The tool design in Honeycomb is relatively simple, streamlining the decision-making process for LLM tool usage, but it may result in the system being unable to obtain more accurate localization information. Honeycomb successfully resolved 38.33% of the issues in the SWE-bench Lite dataset.
Gru (gru, [n. d.]), a workflow-based bug fixing system developed by Gru. It first uses an LLM to select files related to the issue, then the LLM makes decisions on which files to change and how to change them. It generates patches in the format of $(Line_{b},Line_{e},Replace)$, and after the LLM reviews the patches, they are applied to the repository. Gru does not include a reproduction module and successfully resolved 35.67% of the issues in the SWE-bench Lite dataset.
Alibaba Lingma Agent (Ma et al., 2024), a bug fixing system developed by Alibaba that combines code knowledge graphs with LLMs. It constructs a code knowledge graph for the code repository and utilizes LLMs to perform Monte Carlo Tree Search based on issue information to locate code snippets related to the issue throughout the entire repository. This approach effectively alleviates the issue of short context supported by LLMs but places high demands on the reasoning capabilities of the LLMs. It generates patches in Search/Replace format and, for some issues, also validates patches using reproduction scripts. Alibaba Lingma Agent successfully resolved 33.00% of the issues in the SWE-bench Lite dataset.
AutoCodeRover (Zhang et al., 2024), a bug fixing system that combines LLMs with software analysis techniques. It uses Spectrum-based Fault Localization (SBFL) for defect localization, providing functions potentially related to the issue to the LLM, which then generates patches for bug fixing. AutoCodeRover’s localization process offers greater interpretability; however, it requires executing failed unit tests, and its localization effectiveness depends on the quality of the test cases. In practical development, it may struggle to locate defects accurately if tests are missing. AutoCodeRover successfully resolved 30.67% of the issues in the SWE-bench Lite dataset.
Agentless (Xia et al., 2024), a two-stage, workflow-based bug fixing system. It directly uses an LLM to sequentially perform file, function, and line localization. First, it lists the repository’s file structure, allowing the LLM to select the files most relevant to the issue based on the project’s structure. Then, it lists the file framework (including class names, function declarations, etc.) and uses the LLM to identify the functions most relevant to the issue. The model then locates the specific lines that need modification based on the function content. Once the required lines are identified, Agentless provides the context of these lines as input to the model, which generates multiple candidate patches in Search/Replace format. The final patch is selected using a majority voting strategy. Agentless achieves automated bug fixing using a straightforward workflow approach; however, due to the lack of an autonomous decision-making process and missing information in the localization phase, its localization performance is limited. Agentless successfully resolved 27.33% of the issues in the SWE-bench Lite dataset.
Agentless + RepoGraph (Ouyang et al., 2024), a bug fixing system that combines code knowledge graphs with the Agentless approach. It constructs a code knowledge graph for the code repository and provides a graph-based retrieval interface for the LLM, using GRAG (Hu et al., 2024) to retrieve classes/functions related to the issue from the graph. Then, Agentless generates patches based on the localized code snippets, addressing the information loss issue in Agentless’s localization phase. However, it performs poorly on issues that are brief and lack code identifiers, resulting in a gap compared to commercial agent systems. RepoGraph + Agentless successfully resolved 29.67% of the issues in the SWE-bench Lite dataset.
@@
As shown in Table 1, Marscode Agent, which combines three localization strategies and utilizes reproduction for patch validation, achieved the state-of-the-art bug fixing performance. In this paper, we will analyze the localization accuracy of different localization strategies and the impact of reproduction on the final patch generation accuracy, providing guidance for building more reliable bug fixing systems in future work.
Table 1. Leading LLM-based Bug Fixing Systems on SWE-bench Lite.
System Name Type %Resolved Fault Localization Reproduction Patch Generation Format
Using LLMs CodeGraph Software Analysis
MarsCode Agent Commercial 39.33 Search/Replace
Honeycomb 38.33 $\times$ $\times$ ($Line_{b}$, $Line_{e}$, Replace)
Gru 35.67 $\times$ $\times$ $\times$ ($Line_{b}$, $Line_{e}$, Replace)
Alibaba Lingma Agent 33.00 $\times$ Search/Replace
AutoCodeRover Open Source 30.67 $\times$ Search/Replace
Agentless + RepoGraph 29.67 $\times$ $\times$ Search/Replace
Agentless 27.33 $\times$ $\times$ $\times$ Search/Replace
@@ 3. Study Design
In this section, we first introduce the research questions and then introduce the data collection.
@@ 3.1. Research Questions
Motivation: In the SWE-bench Lite leaderboard, the solution rates of various systems vary significantly, and there are substantial differences in the instances that each system can and cannot solve. This discrepancy is usually due to the quality of issue description and the design of the systems themselves. When an issue description is of sufficiently high quality, we expect an LLM-based Agent to be able to resolve it. Therefore, it is necessary to analyze why certain instances with high-quality issue description are not successfully fixed by the Agent, while some instances with low-quality issue description are resolved. Additionally, there are significant differences in the implementation of Agent-based and non-Agent systems, and it is worth investigating the differences in their resolution capabilities.
Approach: We will analyze the differences in the instances solved by various systems, showing how many instances are resolved by all systems and how many instances are not solved by any system. Then, based on the criteria proposed by Agentless[] for evaluating issue descriptions, we will score the quality of the issue descriptions for each instance, where higher scores indicate higher-quality issue descriptions. Subsequently, we will investigate why many high-scoring issues cannot be resolved by any system, while some low-scoring issues can be resolved by all tools. Additionally, we will examine the characteristics of instances that can be resolved by all Agent systems but not by non-Agent systems, as well as those instances that can be resolved by all non-Agent systems but not by Agent systems.
@@
Motivation: Fault localization is a crucial step in bug fixing, as the more accurately the fault is localized, the higher the probability of successfully fixing the bug. Therefore, we need to investigate the differences in fault localization effectiveness among different systems.
@@
Approach: Based on the ground truth, we will compile statistics on the proportion of successfully localized faulty files and the proportion of successfully localized faulty lines for each system in each SWE-bench Lite instance.
RQ3. Effectiveness of Reproduction: How Bug Reproduction in Different Systems Affects Bug Fixing Performance?
Motivation: Bug reproduction is an important step in bug fixing and an essential part of dynamic debugging. Its role is reflected in two aspects. First, the error messages from running the bug reproduction script can be used for fault localization. Second, the bug reproduction script can be used to validate the final generated patch. The higher the quality of the bug reproduction script, the more accurate information it can provide to the Agent, increasing the probability of successfully fixing the bug. Therefore, we need to investigate the impact of bug reproduction on bug fixing.
Approach: We will compile statistics on the adoption rate of the reproduction scripts generated by each system, providing a comparison of the impact of bug reproduction on bug fixing. Additionally, we will analyze the cases that can only be solved with the involvement of reproduction scripts, ans the cases that bug reproduction negatively impact bug fixing.
3.2. Data Collection
In RQ1, we design a scoring system based on the five metrics and corresponding candidate values provided by Agentless, allowing us to evaluate the quality of different issue sets across multiple dimensions. In RQ2, we conduct a reverse analysis of the patches generated by different tools, thereby offering an unbiased evaluation of each tool’s performance in fault localization.In RQ3, to determine the use of reproduction by different systems from their trajectories, we first identify Agentless, RepoGraph+Agentless, and Gru as systems that do not support reproduction, based on keyword matching for "reproduce" and manual analysis. Then, for the remaining four systems, we utilize different heuristic rules to identify the construction of reproduction scripts.

### v2

# LLM-based Agents for Automated Bug Fixing: How Far Are We?

@@ Abstract.
Large language models (LLMs) and LLM-based Agents have been applied to fix bugs automatically, demonstrating the capability in addressing software defects by engaging in development environment interaction, iterative validation and code modification. However, systematic analysis of these agent systems remain limited, particularly regarding performance variations among top-performing ones. In this paper, we examine six repair systems on the SWE-bench Verified benchmark for automated bug fixing. We first assess each system’s overall performance, noting the instances solvable by all or none of these systems, and explore the capabilities of different systems. We also compare fault localization accuracy at file and code symbol levels and evaluate bug reproduction capabilities. Through analysis, we concluded that further optimization is needed in both the LLM capability itself and the design of Agentic flow to improve the effectiveness of the Agent in bug fixing.
copyright: nonecopyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06
@@ 1. Introduction
Large Language Models (LLMs) (Zhao et al., 2023) are advanced machine learning models trained on vast amounts of textual data, capable of understanding and generating human-like text. LLM-based Agents (Xi et al., 2023) are systems that utilize large language models to interact with the environment and accomplish specific tasks. Recently, LLM-based Agents have demonstrated significant influence in automated bug fixing in code repositories (Kang et al., 2023; Zhang et al., 2024; Yang et al., 2024a). Thanks to the powerful natural language processing capabilities of LLMs, these agents can efficiently understand and analyze source code and its associated natural language descriptions, such as user-submitted issue descriptions and code comments. Additionally, through dynamic interaction with local environments (e.g., via terminal), LLM-based agents can retrieve useful information from the code repository, perform code editing and execution, and iteratively validate repair results, thereby improving the accuracy and efficiency of bug fixes. This combination of LLM and environmental feedback has made bug fixing more efficient and feasible than ever before, providing revolutionary new tools for software maintenance and development.
@@
Researchers from both industry (Liu et al., 2024a; Ma et al., 2024; bla, 2025; wan, 2025; eme, 2025; dev, 2025; mid, 2025) and academia (Zhang et al., 2024; Tao et al., 2024; Su et al., 2025) have developed LLM-based agent systems to locate and fix bugs in code repositories. To evaluate the fault localization and repair capabilities of LLMs and various agent systems, Jimenez et.al (Jimenez et al., 2023) proposed the evaluation datasets SWE-bench, with derived versions SWE-bench Lite (a subset of the full benchmark), and SWE-bench Verified (human annotated subset of SWE-bench published recently). These datasets contain real bugs from code repositories and can verify the correctness of the patches generated by Agents through unit tests. Recently, these datasets have become the most influential benchmarks in the field of automated bug fixing, attracting both academic and industrial participants to compete on the corresponding leaderboard [^1] https://www.swebench.com/, with new submissions typically every one or a half week.
@@
However, no work has systematically analyzed the fault localization and repair capabilities of LLM-based Agents or the performance differences among various tools within these Agent systems. Regarding the SWE-bench Verified dataset itself, due to the quality of issue descriptions [^2] An issue description is a concatenated string of ”Issue Titles” and ”Descriptions”. and the complexity of the logical dependencies related to the defects, some instances in the benchmark are easier for Agents to fix, while others are more difficult (Xia et al., 2024). As for the design of the systems, different designs exhibit different planning, reasoning, and problem solving capabilities, i.e. some systems adopting static approaches (gru, 2025) and others adopting dynamic approaches (Zhang et al., 2024). We have also observed significant differences in cases that each system can solve. Therefore, analyzing the solving capabilities of LLM-based Agents on specific instances can not only help us better understand the current performance of Agents but also provide comparative insights for future research directions.
@@
We collected the six systems (i.e. W&B Programmer (wan, 2025), Blackbox AI Agent (bla, 2025), CodeStory Midwit Agent (mid, 2025), Learn-by-interact (Su et al., 2025), Devlo (dev, 2025), Emergent E1 (eme, 2025) with top performances on the SWE-bench Verified Leaderboard and conducted a comprehensive analysis of the performance differences of each system. First, we evaluated the overall performance of LLM-based Agents in bug fixing, including statistics on the instances that all six systems can solve and those that none can solve, and analyzed the reasons behind these results. Next, we investigated the performance differences in fault localization among different systems and their causes, compiling file-level and code symbol-level localization accuracy rates. Finally, we analyzed the impact of bug reproduction on bug fixing.
@@
Through data analysis, we have summarized several insights. To improve bug fixing, it is essential to enhance the model’s reasoning ability, enabling it to accurately identify bug-related information within an issue and reduce noise interference. For multiple potential repair locations, the model should leverage its reasoning capabilities to determine the most relevant location. From the Agentic flow’s perspective, there should be a strong focus on the quality of the issue and attention to multiple suspicious locations in the stack trace. The design should include mechanisms to verify the completeness of patches and consider their global impact. Mechanisms should also be implemented to mitigate the randomness of the model’s output or effectively utilize its diversity. In fault localization, achieving code symbol-level accuracy is more critical than file-level and line-level. During the reproduction process, ensuring the correctness of reproductions is crucial, as incorrect reproductions can result in the failure of the entire solving process.
Contributions: To the best of knowledge, this is the first work to:
@@
  • Study the effectiveness of LLM-based Agents in automatic bug fixing for code repositories.
@@
  • Examine the effectiveness of different LLM-based Agents in Fault Localization and analyze their differences.
  • Investigate the impact of bug reproduction on bug fixing of LLM-based Agents.
@@
  • Summarize the current issues and future research directions for LLM-based Agents in bug fixing.
@@ 2. Background
In this section, we first introduce SWE-bench benchmark and then we introduce the leading LLM-based bug fixing systems.
2.1. SWE-bench Benchmark
@@
After the release of the benchmark, OpenAI’s testing identified several challenging or possibly unsolvable tasks within SWE-bench benchmark, which may cause the benchmark to systematically underestimate the autonomous software engineering capabilities of models. To address these issues, OpenAI collaborated with the authors of SWE-bench to develop a new release of the benchmark that should provide more accurate evaluations, named SWE-bench Verified [^3] https://openai.com/index/introducing-swe-bench-verified/, which contains 500 instances with higher quality. It retains the diversity of SWE-bench but is easier to evaluate. After the release of SWE-bench Verified, it quickly became the most popular benchmark for evaluating large language models and Agent systems on coding performance.
@@ 2.2. Leading LLM-based Bug Fixing Systems
LLM-based Bug Fixing Systems are systems built on Large Language Models (LLMs) that can automatically edit code repositories to fix bugs based on issue reports. Bug fixing is a highly resource-intensive task in software development, requiring developers to reproduce the bugs reported in issue reports, precisely locate defective code snippets within large code repositories, understand the cause of errors, and implement fixes. Automating bug fixing has long attracted widespread attention in both academia and industry. Given the strong logical reasoning and coding capabilities demonstrated by LLMs, numerous works have explored the development of automated bug fixing tools based on LLMs. In this paper, we study six leading LLM-based Bug Fixing Systems, comparing their differences in system design and performance in automated bug fixing, analyzing the shortcomings and limitations of existing systems, and providing direction for future work in building adaptive, high-reliability automated bug fixing systems. Below this section, we will introduce these systems from a technical perspective:
W&B Programmer (wan, 2025), developed by wandb, leverages OpenAI’s o1 model with a high reasoning mode for each step and editing decision. It incorporates a GPT-4o based memory component to compress the agent’s history, a custom Python code editor optimized for model context, and the ability to register "auto-commands" that execute after each edit. To enhance performance, the agent employs five parallel rollouts per instance, followed by a "crosscheck" step using o1 as a tie-breaker to select the best rollout.
Blackbox AI Agent (bla, 2025) developed by Blackbox, is a sophisticated tool designed to enhance productivity and streamline workflows. The agent is equipped with a suite of powerful tools, including get_repo_structure, open_file, create_file, search_file,
and the ability to execute bash commands. These tools enable the agent to efficiently navigate, manipulate, and manage code repositories and files. LLMs leverage these tools in a ReAct (Reasoning and Acting) format to perform a variety of tasks. It can filter and extract relevant content, generate file paths, and validate patches by executing tests. This approach ensures that the agent not only retrieves the necessary information but also verifies the integrity and functionality of the code through rigorous testing.
CodeStory Midwit Agent (mid, 2025), an agent released by CodeStory. They incorporate a suite of specialized tools, including list_files, open_files, str_replace_editor, attempt_completion,
ripgrep_search, and terminal_access for Agent. It utilizes Mon-te Carlo Tree Search(Antoniades et al., 2024) to ensure efficient exploration of the candidate space and filters trajectories by calculating the mean reward of the trajectory’s steps. This approach enhances the performance of the Agent system on SWE-Bench tasks.
Learn-by-interact (Su et al., 2025) employs an agent that directly interacts with code repositories in offline mode to gather extensive high-quality interaction trajectories. It leverages in-context learning to strategically integrate these historical interaction patterns into prompts. This enables the agent to systematically reference and adapt prior successful debugging procedures when addressing novel software defects, resulting in progressively improved bug resolution effectiveness through accumulated experiential knowledge.
Devlo (dev, 2025), an agent released by Devlo. Devlo designs tools such as exploring_repository_structure, creating_file, modifying code, and run_script. Additionally, Devlo employs rigorous testing mechanisms such as reproduction_test and edge_case_test to meticulously filter and validate candidate patches, ensuring robust and reliable outcomes.
Emergent E1 (eme, 2025) initially approaches the problem by engaging multiple expert agents, each equipped with tools such as viewing, creating_file, and str_replace. Subsequently, the final patch is selected from among multiple candidate patches.
@@
In this paper, we will analyze the localization accuracy of different localization strategies and the impact of different components on the final patch generation accuracy, providing guidance for building more reliable bug fixing systems in future work.
@@ 3. Study Design
In this section, we introduce the research questions and the design of issue quality assessment criteria.
@@ 3.1. Research Questions
Motivation: In the SWE-bench Verified leaderboard, resolve rates of different systems vary significantly, with substantial differences in instances each system can and cannot solve, often due to issue description quality and system design. Analyzing why high-quality descriptions sometimes fail while low-quality ones succeed is necessary.
Approach: We will analyze differences in instances solved by various systems, showing how many are resolved by all systems and how many by none. Similar to the assessment done by Agentless on the SWE-bench Lite dataset, we will design an automated method for quality assessment based on state-of-the-art large language models for the SWE-bench Verified dataset. We will investigate why many high-scoring issues are unresolved by any system, while some low-scoring issues are resolved by all systems.
@@
Motivation: Fault localization is crucial in bug fixing since accurate localization increases the probability of fixing the bug. We need to investigate the differences in fault localization effectiveness among different systems.
@@
Approach: Based on the ground truth, we will compile statistics on the proportion of successfully localized faulty files and code symbols (including classes, functions, methods, and top-level statements) for each system in each SWE-bench Verified instance.
RQ3. Effectiveness of Reproduction: How effective is the bug reproduction task currently?
Motivation: Bug reproduction is vital in bug fixing and dynamic debugging. It aids fault localization through error messages and validates patches. Higher-quality reproduction scripts provide more accurate information, increasing the probability of fixing the bug. Investigating the effectiveness of bug reproduction is crucial.
Approach: Due to the difficulty of collecting bug reproduction scripts from various repair systems, we designed an Agent repair system with a mainstream repair process. We collected the content and execution commands of reproduction scripts, subsequently evaluating their quality based on output differences before and after applying the golden patch.
3.2. Issue Quality Assessment Criteria
We design a scoring mechanism based on five metrics, allowing us to evaluate the quality of different issue descriptions across multiple dimensions. Specifically, we drew inspiration from the five metrics used in Agentless (Xia et al., 2024) for evaluating the quality of issue descriptions: Quality of Bug Locations (File/Function/Line), Quality of Reproducible Examples, and Quality of Resolve Solutions. Agentless systematically evaluated the quality of issue descriptions for 300 instances in SWE-bench Lite benchmark using these five metrics. Since SWE-bench Verified was recently proposed, there has been no research conducting similar evaluations on it so far. To fill this gap, we designed clear scoring criteria and utilized the DeepSeek-R1444https://www.deepseek.com/ model, one of the models with the best reasoning abilities and cost-effectiveness, to implement the five-dimensional quality assessment for instances in SWE-bench Verified. One point to note is that the Function-level Fault Localization granularity proposed in Agentless was replaced with a similar but broader granularity of code symbol-level Fault Localization in our study, where code symbols include classes, functions, class methods, and top-level statements. We explain the condensed scoring criteria for the five dimensions as follows, and full prompts for the LLM to make the quality assessments can be obtained in our homepage [^5] https://github.com/ResearchOpenRepos/bug_fixing_agent_empirical_study/tree/master/issue_quality_annotation.
Quality of File-level Fault Localization: We identify edited code files based on the golden patch (human-written patch, considered ground truth) and treat them as golden files. We then input the issue description and the corresponding golden files to the LLM to determine the extent to which golden files appear in issue descriptions. The scoring criteria are:
  1. (1)
    StackTrace (3). The issue description contains stack trace information, which contains the modified file name.
  2. (2)
    Keyword (2). The issue description contains the file name and its path prefix. The path prefix is required to uniquely identify the file entity. For example, for path/to/file_reader.py, either path/to/file_reader.py or path/to/file_reader is acceptable. File path and name may appear in different locations but both must be present in the issue description.
  3. (3)
    Natural Language (1). The issue description provides the file name in natural language without path prefix. The file name cannot be split - file_reader.py or file_reader is valid, but file and reader appearing separately does not qualify.
  4. (4)
    No Information (0). This applies when keywords appear separately without forming the complete file name, for example, file and reader appear individually rather than as file_reader, or when no relevant keywords appear.
Since a single case may contain multiple golden files, the LLM is required to evaluate each file based on the aforementioned scoring criteria, with the maximum score ultimately being selected as the annotation result for that case.
[…]
  1. (1)
    StackTrace (3). The issue description contains stack trace information, which contains the target code symbol name.
  2. (2)
[…]
  3. (3)
    Natural Language (1). The issue description mentions the code symbol name in natural language without file path. Following the same principle as file-level localization, these symbol names must remain intact as atomic units, e.g., the symbol get_name would not be considered matched if only its constituent parts (get or name) appear independently.
  4. (4)
    No Information (0). The issue description contains neither direct nor sufficiently indicative references to the target symbol. Partial or indirect mentions (e.g., separate components like get and name appearing without the complete symbol get_name) are treated as no symbol information.
Quality of Line-level Fault Localization: We use an out-of-the-box tool named delta [^6] https://github.com/dandavison/delta to convert the standard diff format golden patch into a show-line golden patch, displaying line numbers before and after the patch at the beginning of each line. This annotation aids the model in understanding code modifications and accurately matching line numbers. We then provide the issue description and the show-line golden patch to the LLM. The scoring criteria are:
  1. (1)
    StackTrace (3): The description must contain a stack trace that explicitly includes at least one pair of <file_name, line_number> corresponding to the modified code locations in the patch.
  2. (2)
    Keyword (2): At least one pair of <file_name, line_number> appears in the issue text, but not appears in a stack trace.
  3. (3)
    Natural Language (1): The issue description contains a natural language explanation of the buggy line(s), but does not specify the exact line number(s). For example, if the patch contains result = get_content(url, proxy, timeout), the description including result = get_content(url, proxy, timeout) or get_content(url, proxy, timeout) qualifies. Isolated identifiers like proxy or timeout do not qualify.
  4. (4)
    No Information (0): No useful line-related information appears in the issue description.
Quality of Reproducible Examples: We provide only the issue description, without additional ground truth information. Consequently, our approach only requires the LLM to detect the presence of reproducible examples, without evaluating their semantic correctness. The scoring criteria are:
  1. (1)
    Contains REs (3). The issue description includes complete, directly runnable code to reproduce the problem without additional setup.
  2. (2)
    Contains Partial REs (2). The description provides partial code, containing the main part needed but lacking some information, requiring extra context or setup to run.
  3. (3)
    Info in NL (1). The description provides a detailed, step-by-step explanation in natural language on how to reproduce the problem, without code snippets.
  4. (4)
    Not Enough Info (0). There is insufficient information to reproduce the problem.
Quality of Resolve Solutions: The model receives the issue description and the golden patch. Scoring criteria are:
  1. (1)
    Exact Patch (3). The issue description provides a code solution that is exactly the same as the correct patch. If there is any semantic discrepancy or if the code snippet only covers part of the standard patch rather than the entire content, this option should not be selected.
  2. (2)
    Complete Steps in NL (2). The issue description provides a complete natural language solution that is semantically consistent with the correct patch. One could reasonably write the correct patch based on this description without additional guidance. If the description only covers part of the standard patch, this option should not be selected.
  3. (3)
    Some Steps in NL (1). The issue description provides a partial natural language solution that is not semantically complete compared to the provided patch. One could write part of the correct patch based on this description but not all. This occurs when users have incomplete understanding and provide partial solutions.
  4. (4)
    No Solution (0). The issue description does not provide any solution description - the user only points out the problem without suggesting how to fix it.
  5. (5)
    Misleading Information (-1). The issue description includes the user’s guess on how to solve the problem, but this guess is completely wrong and off track compared to the correct patch, making it misleading information.
3.3. Reliability of the Scoring System
To validate the reliability of the scoring mechanism using DeepSeek R1, we performed a cross-validation against deterministic regex rules designed to identify structured cues such as stack traces in issue descriptions. Specifically, we analyzed all 500 issues and found disagreements in 34 cases. A detailed inspection revealed that DeepSeek substantially outperforms regex-based methods. Regex rules, though precise, often fail to generalize across diverse stack trace formats, which vary by Python version, bug-tracking tool, and runtime environment (e.g., Jupyter Notebook). Furthermore, many issue reports describe stack traces or error locations using natural language rather than standardized formatting. In such cases, DeepSeek’s semantic understanding allows it to successfully identify valuable debugging signals that regex rules miss. This result supports the use of DeepSeek R1 as a more robust and adaptable scoring mechanism for assessing issue quality.